# Statistical Taylor Expansion: A New and Path-Independent Method for Uncertainty Analysis 111

### v16

@@ 2.11 Dependency Tracing
$\delta^{2}(fg)=$ $\;\overline{f^{2}g^{2}}-\overline{fg}^{2};$ (2.55)
$\delta^{2}f(g)=$ $\;\overline{f(g)^{2}}-\overline{f(g)}^{2};$ (2.56)
@@
  • Formula (2.54) expresses $\delta^{2}(f+g)$, whose dependency tracing is illustrated by $\delta^{2}(f-f)=0$, and $\delta^{2}(f(x)+g(y))=\delta^{2}f+\delta^{2}g$, with the latter corresponding to Formula (2.15). Formula (LABEL:eqn:_Taylor_inner_dependency) shows that Formula (2.3) applies Formula (2.54) between any two terms in the Taylor expansion in Formula (2.8).
  • Formula (2.55) expresses $\delta^{2}(fg)$, illustrated by $\delta^{2}(f/f)=0$, $\delta^{2}(ff)=\delta^{2}f^{2}$, and $\delta^{2}(f(x)g(y))=\overline{f}^{2}(\delta^{2}g)+(\delta^{2}f)\overline{g}^{2}+(\delta^{2}f)(\delta^{2}g)$, with the latter corresponding to Formula (2.17).
  • Formula (2.56) shows $\delta^{2}f(g(x))$, whose dependency tracing is demonstrated by $\delta^{2}(f^{-1}(f))=(\delta x)^{2}$.
  • Formula (2.57) gives the variance of the linear transformation of a function, which can be applied to Formula (2.54) and (2.55) for more general dependency tracing.
Variance arithmetic employs dependency tracing to ensure that the calculated mean and variance satisfy statistics rigorously. However, dependency tracing comes at a cost: variance calculations are generally more complex than value calculations and exhibits a narrower convergence range for input variables. Dependency tracing also implies that the results of statistical Taylor expansion must remain path independent.
@@ 2.12 Traditional Execution and Dependency Problem
  • Furthermore, traditional approaches frequently apply approximations to result values during execution. Under the statistical Taylor expansion, Formula (2.3) shows that the variance converges more slowly than value in statistical Taylor expansion. Consequently, approximation strategies should prioritize variances than values. Section 7 illustrates this principle through a first-order approximation used in computing a matrix determinant.
  • Traditionally, results from mathematical library functions are accepted without scrutiny, with accuracy assumed down to the last bit. As demonstrated in Section 9, statistical Taylor expansion enables the detection of numerical errors within these functions and requires that they be recalculated with uncertainty explicitly incorporated into the output.
  • In conventional practice, an analytic expression is often decomposed into simpler, ostensibly and independent arithmetic operations such as negation, addition, multiplication, division, square root, and library calls. However, this decomposition introduces dependency problems in floating-point arithmetic, interval arithmetic, and statistical Taylor expansion. For example, if $x^{2}-x$ is calculated as $x^{2}-x$, x(x1)x(x-1), and $(x-\frac{1}{2})^{2}-\frac{1}{4}$, only $(x-\frac{1}{2})^{2}-\frac{1}{4}$ gives the correct result, while the other two give wrong results for wrong independence assumptions between $x^{2}$ and $x$, or between x1x-1 and $x$, respectively.
  • Similarly, large calculations are often divided into sequential steps, such as computing $f(g(x))$ as $f(y)|_{y=g(x)}$. This approach fails in statistical Taylor expansion because dependency tracing within $g(x)$ affects $f(g(x))$. In this context, $\overline{f(g(x))}\neq\overline{f(y)|_{y=g(x)}}$ and $\delta^{2}f(g(x))\neq\delta^{2}f(y)|_{y=g(x)}$. The path dependence of $f(y)|_{y=g(x)}$ and $\delta^{2}f(y)|_{y=g(x)}$ are evident in cases such as $\overline{(\sqrt{x})^{2}}>\overline{\sqrt{x^{2}}}$ and $\delta^{2}(\sqrt{x})^{2}>\delta^{2}\sqrt{x^{2}}$.
@@ 3 Variance Arithmetic
Variance arithmetic implements statistical Taylor expansion. Because of the finite precision and limited range of conventional floating-point representation, $\zeta(n)$ can only be computed to limited terms. Consequently, the following numerical rules are introduced:
@@
  • monotonic: As a necessary condition for convergence, the last $20$ terms of the expansion must decrease monotonically in absolute value, ensuring that the probability of the expansion exhibiting an absolute increase is no more than $2^{-20}=9.53\;10^{-7}$. Unless all the remaining terms in the expansion are known to be precisely zeros, each expansion is executed to the full 448 terms for the monotonicity check.
  • positive: At every order, the expansion variance must be positive.
  • stable: To avoid truncation error [12], the value of the last expansion term must be less than $5.73\;10^{-7}$ times of both the result uncertainty and the result absolute value, in which $5.73\;10^{-7}$ is the bounding leakage for Gaussian distribution with bounding range $\kappa=5$. This rule ensures sufficiently fast convergence in the context of monotonic convergence.
  • reliable: At every order, the uncertainty of the variance must be less than $1/5$ times of the value of the variance.
For simplicity of discussion, This paper confines the calculation of variances to ideal variances $\widehat{\delta^{2}}f$ and assumes that the input distribution is Gaussian with $\kappa_{s}=5$. Furthermore,the Taylor coefficients in Formula (2.8) and (2.11) are assumed to be precise.
3.1 Numerical Representation
Variance arithmetic represents an imprecise value $x\pm\delta x$ using a pair of 64-bit standard floating-point numbers. All other conventional numerical numbers must be converted to this format.
If the least $20$ significant bits of the significand in a standard floating-point number are all 0, the value is considered precise, representing a 2’s fractional with a probability no less than $1-2^{-20}=1-2.384\;10^{-7}$. Otherwise, a standard floating-point value is considered imprecise with its uncertainty defined as $1/\sqrt{3}$ times of the ULP of the value, where ULP refers to the Unit in the Last Place in conventional floating-point representation [8]. This follows from the fact that the pure rounding error in round-to-nearest mode is uniformly distributed within half bit of the significand of a floating-point value [29].
If an integer number is within the range $2^{53}+1,+2^{53}-1]$ of the significand of a 64-bit standard floating-point number, its uncertainty is zero. Otherwise, it is first converted to a conventional floating-point value before being transformed into an imprecise value.
Variance arithmetic uses floating-point arithmetic for computation.
3.2 Finite
For $(1\pm\delta x)^{-2}$, the Taylor coefficient increases with the expansion order $n$ as $(-1)^{n}(n+1)$. When $\delta x=0.5$, this growth causes the result variance to diverge to infinity.
@@ 3.3 Monotonic
Figure 11: Measured upper bound $\delta x$ (left y-axis) for $(1\pm\delta x)^{c}$ across different values of $c$ (x-axis) for uniform uncertainty. The corresponding uncertainty bias and uncertainty are also shown (right y-axis). The x-axis is expressed in units of $\pi$. When $c$ is a natural number, $\delta x$ has no upper bound; however, such cases are omitted in the figure.
From Formula (2.8), the convergence condition applied to $\log(x\pm\delta x)$ is $P(x)\leq 1/\kappa=1/5$, which is numerically confirmed as $P(x)\lesssim 0.20086$. Beyond this upper bound, the expansion is no longer monotonic. Variance arithmetic rejects the distributional zero of $\log(x)$ in the range of $x-\delta x,x+\delta x]$ statistically due to the divergence of Formula (2.23) mathematically, with $\zeta(2n)$ providing the connection between these two perspectives.
For $e^{x\pm\delta x}$ the convergence holds for $\delta x\lesssim 19.864$ regardless of $x$, while the result $\delta\log(x\pm\delta x)\lesssim 0.213$ regardless of $x$. These limits follow directly from the relationship $\delta x\rightarrow P(e^{x})$ and $P(x)\rightarrow\delta\log(x)$, as indicated in Formula (2.20) and (2.23).
From Formula (2.49), and except when $c$ is a natural number, Formula (2.29) for $(x\pm\delta x)^{c}$ converges near $P(x)\simeq 1/\kappa=1/5$, with the upper bound $P(x)$ increasing with $c$. This trend is approximately confirmed in Figure 10, and the expansion is no longer monotonic beyond the upper bound $P(x)$. Qualitatively, $\delta^{2}1/x$ converges more slowly than $\delta^{2}\sqrt{x}$, consistent with Formula (2.36) and (2.35).
When the input uncertainty is Uniformly distributed, Figure 11 illustrates that the input uncertainty upper bound is close to $1/\kappa=/\sqrt{3}$ according to Formula (2.49). The bounded momentum $\zeta(2n)$ for Uniform distribution is much less than that for Normal distribution, such that the expansion contains $652$ terms instead of $448$ terms. The upper bounds of input uncertainty in Figure 10 and 11 looks identical, however, the resulting uncertainty biases and uncertainties in Figure 10 are $3$ order-of-magnitude smaller than those in Figure 11 when the exponent $c$ is less than 0. Even a result converges, it can still be abandoned due to too large uncertainty bias or uncertainty.
@@ 3.4 Positive
In some cases, the variance expansion may yield negative results, as in Formula (2.26) for $sin(x\pm\delta x)$. Figure 12 shows that the upper bound of $\delta x$ for $sin(x\pm\delta x)$ varies periodically between $0.318\pi$ and $0.416\pi$. Beyond this upper bound, the expansion is no longer positive.
In Figure 13, when the bounded momentum $\zeta(2n)$ for Uniform distribution is used instead of that for Normal distribution, the input uncertainty upper bound is $4$ times larger than that in Figure 12 but is still periodic. The resulting uncertainty increases not much and still remains less than $1$.
3.5 Stable
The unstable condition happens rarely.
3.6 Reliable
The condition of not being reliable seldom happens.
@@ 4.1 Verification Methods and Standards
  • Value deviation: the standard deviation of the value errors.
  • Normalized error: the ratio of a value error to the corresponding uncertainty.
@@
  • Uncertainty mean: the sample mean of the result uncertainties.
  • Uncertainty response: the relationship between input uncertainties and output uncertainties.
  • Calculation response: the relationship between the amount of calculation and output uncertainties.
One objective of uncertainty-based computation is to account precisely for all input errors from every source, thereby achieving ideal coverage:
  1. 1.
    The error deviation is exactly $1$.
  2. 2.
    The error distribution should follow Normal distribution when an imprecise value is expected, or Delta distribution when a precise value is expected.
If the precise result is unknown, the resulting normalized error distribution can be used to assess whether ideal coverage is achieved.
@@
When an input contains unspecified errors, such as numerical errors in library functions, Gaussian noise with progressively increasing deviations can be added, until ideal coverage is attained. The minimal noise deviation required provides a good estimate of the magnitude of the unspecified input uncertainties. Achieving ideal coverage serves as a necessary verification step to ensure that Formula (2.3) or (2.3) have been applied correctly within the given context. The input noise range that yields ideal coverage defines the ideal application range for input uncertainties.
Besides coverage, other validation criteria are also used:
  • The uncertainty response should match the expected functional form. For a linear function, the output uncertainties should increase linearly with the input uncertainties, with the ratio of output to input uncertainty means defined as uncertainty response ratio.
  • The calculation response should follow the expected trend; for instance, increasing the number of calculations should result in larger output uncertainties.
Noises of varying deviations can also be added to the input to evaluate a function’s uncertainty response.
4.2 Types of Uncertainties
There are five primary sources of result uncertainty in a calculation [1][12][29]:
  • Input uncertainties: The examples presented in this paper demonstrate that when the precision of input uncertainties is $10^{-15}$ or larger, variance arithmetic can achieve ideal coverage for input uncertainties.
  • Rounding errors: Empirical results indicate that variance arithmetic provides proper coverage for rounding errors.
  • Truncation errors: Variance arithmetic avoids truncation errors with its stable rule. However, using Formula (5.1) for polynomial expansion can result in truncation errors when the choice of the expansion order is insufficient, as illustrated in Figure 14 .
  • External errors: External errors are value errors not specified in the input uncertainties, such as numerical errors in library functions. Section 9 examines the effects of numerical errors of library sine function, showing that when these external errors are sufficiently large, neither ideal coverage nor proper coverage can be achieved. This finding indicates that library functions must be recalculated to explicitly include the corresponding uncertainty for each computed value.
  • Modeling errors: Modeling errors arise when an approximate analytic solution is used, or when a real-world problem is simplified to make a solution tractable. For example, Section 9 demonstrates that the discrete Fourier transform (DFT) is only an approximation of the mathematically defined continuous Fourier transform (FT), and therefore contains modeling errors. Conceptually, modeling errors originate in mathematics within and are thus outside the domain of statistical Taylor expansion.
4.3 Types of Calculations to Verify
Algorithms of distinct natures, with each representative of its respective category, are required to test the broad applicability of variance arithmetic [29]. An algorithm can be categorized by comparing the amount of its input and output data as [29]:
  • Application,
  • Transformation,
  • Generation,
  • Reduction.
An application algorithm computes numerical values from an analytic formula. Through statistical Taylor expansion, variance arithmetic applies directly to analytic problems. For example, for the catastrophic cancellation example of Formula (1.2), Formula (4.3) shows that variance arithmetic bound the rounding error of $1$ with uncertainty $1$:
$64919224$ $\times 205117922-159018721\times 83739041=$
$13316075197586562.-13316075197586560.\pm 1=2.\pm 1;$ (4.1)
A transformation algorithm produces output data of approximately the same quantity as its input, with the overall information content remaining largely unchanged. For reversible transformations, a unique requirement is to recover every original input for both value and uncertainty after a roundtrip transformation that is performing a forward transformation followed by its reverse transformation. The discrete Fourier transform (DFT) is a typical reversible transformation algorithm: it has the same amount of input and output data, and its output can be transformed back into the input using essentially the same process. A test of variance arithmetic using the fast Fourier transform (FFT, which is an implementation of DFT) algorithms is presented in Section 9.
A generation algorithm produces substantially more output data than input data. Such algorithms encode mathematical knowledge into data. Certain generation algorithms are purely theoretical calculations that involve no imprecise input, so all resulting uncertainty arises solely from rounding errors. Section 10 presents a generation algorithm that generates a sine function table using trigonometric relations and two precise inputs: $sin(0)=0$ and $sin(\pi/2)=1$.
A reduction algorithm yields significantly fewer output data than input data, as in numerical integration, or in the statistical characterization of a data set. In this process, certain information is lost while other information is extracted. As a statistical approach, variance arithmetic inherently carries out statistical reduction. For example, the result of averaging $N$ imprecise values with each uncertainty close to $\delta x$ is $\overline{x}\pm\frac{\delta x}{\sqrt{N}}$, reflecting central limit theorem [4] naturally.
@@ 5 Polynomial
5.1 Tracking Rounding Error
@@
Variance arithmetic can track rounding errors effectively without the need for additional rules.
Figure 14 shows the residual error of $\sum_{j=0}^{224}x^{j}-\frac{1}{1-x}$, where the polynomial $\sum_{j=0}^{224}x^{j}$ is computed using Formula (5.1), $\frac{1}{1-x}$ is computed using Formula (2.27), and $x$ is initiated as a floating-point value. Because $\eta(2n)$ is limited to $2n\leq 448$, Formula (5.1) for polynomial evaluation is restricted to $N\leq 224$, so that $\sum_{j=0}^{224}x^{j}$ has lower expansion order than that of the statistical Taylor expansion of $\frac{1}{1-x}$. Figure 14 shows:
  • When $x\in[-0.73,0.75]$, the required expansion order is no more than $224$, which indicates that the residual error reflects solely the rounding error between $\sum_{j=0}^{224}x^{j}$ and $\frac{1}{1-x}$. A detailed analysis indicates that the maximal residual error is $4$ times the ULP of $\frac{1}{1-x}$. The calculated uncertainty bounds the residual error effectively for all $x\in[-0.73,0.75]$.
  • When $x\not\in[-0.74,+0.75]$, the required expansion order exceeds $224$, so that the residual error arises from the insufficient expansion order of $\sum_{j=0}^{224}x^{j}$. The residual error magnitude increases as $|x|\rightarrow 1$, reaching approximately $50$ when $x=0.98$.

### v17

@@ 2.8 Dependency Tracing
  • Formula (2.32) expresses $\delta^{2}(f+g)$, The dependency tracing of is illustrated by $\delta^{2}(f-f)=0$, and $\delta^{2}(f(x)+g(y))=\delta^{2}f+\delta^{2}g$, with the latter corresponding to Formula (2.11).
  • The dependence tracing of $\delta^{2}(fg)$ is illustrated by $\delta^{2}(f/f)=0$, $\delta^{2}(ff)=\delta^{2}f^{2}$, and $\delta^{2}(f(x)g(y))=\overline{f}^{2}(\delta^{2}g)+(\delta^{2}f)\overline{g}^{2}+(\delta^{2}f)(\delta^{2}g)$, with the latter corresponding to Formula (2.13).
  • The dependency tracing of $\delta^{2}f(g(x))$ is demonstrated by $\delta^{2}(f^{-1}(f(x)))=(\delta x)^{2}$. For a reversible transformation such as matrix reversion or FFT, after a round-trip transformation which is a forward transformation followed by a reverse transformation, the original input should be restored.
  • Formula (2.33) gives the variance of the linear transformation of a function, which can be applied to other formulas for more general dependency tracing.
Statistical Taylor expansion employs dependency tracing to ensure that the calculated mean and variance satisfy statistics rigorously. Dependency tracing also implies that the results of statistical Taylor expansion must remain path independent. However, dependency tracing comes at a cost: variance calculations are generally more complex than value calculations and exhibits a narrower convergence range for input variables.
@@ 2.9 Traditional Execution and Dependency Problem
  • Furthermore, traditional approaches frequently apply approximations to result values during execution. Under the statistical Taylor expansion, Formula (2.3) shows that the variance converges more slowly than value in statistical Taylor expansion. Consequently, approximation strategies should prioritize variances than values.
  • Traditionally, results from mathematical library functions are accepted without scrutiny, with accuracy assumed down to the last digit. As demonstrated in Section 10, statistical Taylor expansion enables the detection of numerical errors within these functions and requires that they be recalculated with uncertainty deviations explicitly incorporated into the output.
  • In conventional practice, an analytic expression is often decomposed into simpler, ostensibly and independent arithmetic operations such as negation, addition, multiplication, division, square root, and library calls. However, this decomposition introduces dependency problems. For example, if $x^{2}-x$ is calculated as $x^{2}-x$, x(x1)x(x-1), and $(x-\frac{1}{2})^{2}-\frac{1}{4}$, only $(x-\frac{1}{2})^{2}-\frac{1}{4}$ gives the correct result, while the other two give wrong results for wrong independence assumptions between $x^{2}$ and $x$, or between x1x-1 and $x$, respectively.
  • Similarly, large calculations are often divided into sequential steps, such as computing $f(g(x))$ as $f(y)|_{y=g(x)}$. This approach also introduces the dependency problem by ignoring dependency tracing within $g(x)$ affecting $f(g(x))$, such as $\overline{(\sqrt{x})^{2}}>\overline{\sqrt{x^{2}}}$ and $\delta^{2}(\sqrt{x})^{2}>\delta^{2}\sqrt{x^{2}}$.
@@ 3 Variance Arithmetic
Variance arithmetic implements statistical Taylor expansion. It represents an imprecise value $x\pm\delta x$ using a pair of 64-bit standard floating-point numbers and uses floating-point arithmetic for computation.
Because of the finite precision and limited range of conventional floating-point representation, $\zeta(n)$ can only be computed to limited terms. Consequently, the following numerical rules are introduced:
@@
  • monotonic: As a necessary condition for convergence, the last $20$ terms of the expansion must decrease monotonically in absolute value, ensuring that the probability of the expansion exhibiting an absolute increase is no more than $2^{-20}\simeq 9.53\;10^{-7}$.
  • stable: To avoid truncation error [5], the absolute value of the last expansion term must be less than $\epsilon$ times of both the result deviation and the result absolute value, in which $\epsilon\simeq 5.73\;10^{-7}$ is the bounding leakage for Gaussian distribution with $\hat{\kappa}=5$. This rule ensures sufficiently fast convergence in the context of monotonic convergence.
  • positive: At every expansion order, the expansion variance must be positive.
  • reliable: At every order, the deviation of the variance must be less than $1/5$ times the value of the variance.
For simplicity of discussion, the Taylor coefficients in Formula (2.4) and (2.7) are assumed to be precise.
@@ 3.1 Monotonic
Figure 9: Measured upper bound $\delta x$ (left y-axis) for $(1\pm\delta x)^{c}$ across different values of $c$ (x-axis) for uniform uncertainty. The corresponding uncertainty bias and uncertainty are also shown (right y-axis). The x-axis is expressed in units of $\pi$. When $c$ is a natural number, $\delta x$ has no upper bound; however, such cases are omitted in the figure. When $c\in[0,1]$, uncertainty bias is negative so it does not show up in the plot.
Beyond an upper bound $\delta x$, the expansion is no longer monotonic for $e^{x\pm\delta x}$, $\log(x\pm\delta x)$, and $(x\pm\delta x)^{c}$.
For Gaussian input uncertainty with $\hat{\kappa}=5$:
  • For $e^{x\pm\delta x}$, $\delta x\lesssim 19.864$ and $P(e^{x\pm\delta x})\lesssim 1681.767$ regardless of $x$. These limits follow directly from the relationship $\delta x\rightarrow P(e^{x})$, as indicated in Formula (2.15).
  • For $\log(x\pm\delta x)$, $P(x)\lesssim 0.20086$ and $\delta\log(x\pm\delta x)\lesssim 0.213$ regardless of $x$. These limits follow directly from the relationship $P(x)\rightarrow\delta\log(x)$, as indicated in Formula (2.17).
  • For $(x\pm\delta x)^{c}$, except when $c$ is a natural number, the upper bound $P(x)$ is close to $1/5$ but increasing with $c$. This trend is displayed in Figure 8.
Similar trends holds when the input uncertainty is uniform but with different values. For example, after the respective upper bound $P(x)$ for $(x\pm\delta x)^{c}$ are normalized by the corresponding $\hat{\kappa}$, they show almost the same trends of increasing with $c$. Figure 9 shows the upper bound $P(x)$ for $(x\pm\delta x)^{c}$.
@@ 3.2 Positive
Besides convergence, the variance expansion may yield negative results, as in Formula (2.20) for $sin(x\pm\delta x)$. Figure 10 shows that the upper bound of $\delta x$ for $sin(x\pm\delta x)$ varies periodically between $0.318\pi$ and $0.416\pi$ for Gaussian input uncertainty. Beyond this upper bound, the result deviation is no longer positive at some expansion order. Conceptually, the upper-bound for $\delta x$ should not be much more than $\pi/2$ to make the result meaningful, which Figure 10 satisfies. Similar trend holds when the input uncertainty is uniform but with larger upper bound $\delta x$, as shown in Figure 11.
3.3 Floating-Point Rounding Errors
Variance arithmetic incorporates floating-point rounding errors as the $\delta x$ when converting a floating-point value $x$ into $x\pm\delta x$. Unless all the least 20 bits of the significand of $x$ are zero, $\delta x$ is assumed to be $1/\sqrt{3}$ times of the ULP of $x$, where ULP refers to the Unit in the Last Place in conventional floating-point representation [8], because rounding errors are shown to be uniformly distributed within ULP [17].
@@ 4 Verification of Variance Arithmetic
  • Normalized error: the ratio of a value error to the corresponding result deviation from statistical Taylor expansion.
@@
Once input error from every source is accounted precisely, ideal coverage is achieved in either contexts:
  • Distribution Test: When comparing the calculated mean and deviation with the result data set, the error deviation is exactly $1$ and the error distribution is Normal. Empirically, error distributions are close to Normal when error deviation are close to 1, regardless of input uncertainty distribution, and such convergence occurs rapidly [17]. This is due to central limit theorem [4].
  • Value Test: When comparing the calculated value and the result value in a corresponding data set, the error deviation is much less than $1$ and the error distribution is Delta. For example, a round-trip test is a value test.
@@
When an input contains unspecified errors, such as numerical errors in library functions or floating-point rounding errors, Gaussian noise with progressively increasing deviations can be added, until ideal coverage is attained. The minimal noise deviation required provides a good estimate of the magnitude of the unspecified input uncertainty deviations. Achieving ideal coverage serves as a necessary verification step to ensure that statistical Taylor expansion has been applied correctly within the given context. The input noise range that yields ideal coverage defines the ideal application range for the analytic function.
@@ 5 Polynomial
Because the maximal expansion term using Formula (5.1) is $\tilde{x}^{2N}$, $N$ in Formula (5.1) can only reach half of the maximal expansion order of Formula (2.3), for example, $224=448/2$ when the input uncertainty is assumed to be Gaussian.
5.1 Residual Error
@@
Figure 12 shows the residual error of $\sum_{j=0}^{224}x^{j}-1(1-x)$. It also displays the required expansion orders for 1/(1x)1/(1-x), which are all less than $224$. Therefore, the residual error reflects solely the rounding error between $\sum_{j=0}^{224}x^{j}$ and $\frac{1}{1-x}$. A detailed analysis indicates that the maximal residual error is $4$ times the ULP of 1/(1x)1/(1-x). In all cases, the calculated uncertainty bounds the residual error effectively for all $x$, with an error deviation of $2.60$ when the expansion order is less than $224$. Variance arithmetic can provide proper coverage for rounding errors.